Using shell method, find the volume of R, when it is bounded by √x + √y = √a , x =

0 , y = 0 about the line x = a.


1
Expert's answer
2022-05-30T23:01:40-0400


"\sqrt{x}+\sqrt{y}=\sqrt{a},"

"x=(\sqrt{a}-\sqrt{y})^2=a-2\sqrt{a}\sqrt{y}+y"




"V=\displaystyle\int_{0}^a2\pi y(a-(a-2\sqrt{a}\sqrt{y}+y))dy"

"=2\pi\displaystyle\int_{0}^a(2\sqrt{a}y^{3/2}-y^2)dy"

"=2\pi[\dfrac{4\sqrt{a}y^{5/2}}{5}-\dfrac{y^3}{3}]\begin{matrix}
 a \\
 0
\end{matrix}"

"=\dfrac{14\pi a^3}{15} ({units}^3)"

"\dfrac{14\pi a^3}{15}" cubic units.